### v1

@@ 11. A lci-assignment
Let $\mathbf{Q}$=“$(S_{2})$” and $\mathbf{P}$=“lci.” Let $A\in\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$. Intuitively, we want to define $\mathfrak{c}(A)$ to be the Fitting invariant of modules $C_{n}(A/R)$ showing up in [Briggs-Iyenger-Cotangent-Complex], where $R$ is a regular local ring mapping surjectively to $A$. The flatness of $C_{n}(A/R)$ characterizes lci. However, these modules depend on the choice of $R$ and a projective resolution (in a way that does not change the Fitting invariant, however), and are fragile along ascent (i.e. still involve non-finite modules). We will work with $C_{n}(A/\mathbf{Z})$ instead, which gives the same Fitting invariant. We use standard notations for derived categories, and cohomological conventions for cotangent complexes, as in [stacks].
Discussion 11.1.
@@
The collection of all such $C^{a}(P^{\bullet})$ is denoted $\mathcal{C}^{a}(L)$. For $X,Y\in\mathcal{C}^{a}(L)$, there exist projective modules $P,Q$ so that $X\oplus P\cong Y\oplus Q$, see [Briggs-Iyenger-Cotangent-Complex, (7.2)]. We write $C^{a}(L)$ for an unspecified element in $\mathcal{C}^{a}(L)$. The flat and projective dimensions of $C^{a}(L)$ are well-defined.
If $L$ has tor-amplitude in $[a,b]$, then $C^{a}(L)$ is flat. Indeed, let $P^{\bullet}$ represent $L$, then $\tau_{\geq a}(P^{\bullet}\otimes_{A}M)$ represents $L\otimes_{A}^{L}M$ for all $A$-modules $M$, as $L\otimes_{A}^{L}M\in D^{[a,b]}(A)$. Unwinding the definitions, we see $C^{a}(P^{\bullet})\otimes_{R}M=C^{a}(P^{\bullet})\otimes^{L}_{R}M$, as desired. This also tells us if $C^{a}(L)$ has projective dimension $p<\infty$, then $L$ has projective-amplitude in $[a-p,b]$.
@@
[…]
@@
Apply Discussion 11.2 to the triangle
[…]
@@
[…]
where $P=C^{a+1}(L_{R/\mathbf{Z}}\otimes^{L}_{R}A)$, $H=H^{a}(L_{R/\mathbf{Z}}\otimes^{L}_{R}A)$. Since $R$ is lci, $L_{R/\mathbf{Z}}$ has tor-amplitude in $[-1,0]$ [avramov-ci, (1.2) and (5.1)], so $H=0$. Moreover, every flat $A$-module has projective dimension $\leq\dim A$ [Raynaud, Seconde partie, Corollaire 3.2.7]. Therefore $L_{R/\mathbf{Z}}$ has projective-amplitude in $[-\dim A-1,0]$, so $P$ is projective, and we get $C^{a}(L_{A/\mathbf{Z}})\oplus P\cong C^{a}(L_{A/R})$. It remains to observe $C^{a}(L_{A/R})$ is finite up to projective summands, as $L_{A/R}\in D_{Coh}(A)$ [stacks, Tag 08PZ]. ∎
@@
If $A$ is countable, then we can improve $-\dim A-1$ to $-2$, see [Raynaud, Seconde partie, Corollaire 3.3.2]. We could, if necessary, work extensively with countable rings, via a Löwenheim–Skolem type argument, cf. [Lyu-elementary-subring].
We would like to define the Fitting invariant of $C^{a}(L_{A/\mathbf{Z}})$ to be that of $M$; we will show this is well-defined. Before that, note the following variant of the main theorem of [Briggs-Iyenger-Cotangent-Complex].
@@
Take the same exact sequence as in the proof of Lemma 11.3. We have $H=0$ and $P$ is flat. Thus $C^{a}(L_{A/\mathbf{Z}})$ is flat if and only $C^{a}(L_{A/R})$ is flat, if and only if $R\to A$ is lci ([Briggs-Iyenger-Cotangent-Complex, Theorem B] and [avramov-ci, (1.2)]), if and only if $A$ is lci [avramov-ci, (5.4) and (5.9)]. ∎
@@ 2. $(ii)$
We have $X=\operatorname{colim}_{\alpha}X_{\alpha}$ as filtered colimits commute with finite limits. Therefore the inclusion $M^{\prime}\subseteq X$ factors through some $X_{\alpha}$ [stacks, Tag 0G8P]. By [stacks, Tag 058M] $M^{\prime}\subseteq X$ is pure (i.e. universally injective), so $M^{\prime}\to X_{\alpha}$ is also pure, thus split [stacks, Tag 058L]. On the other hand $X_{\alpha}\cong M\oplus L_{\alpha}$ as $A$-modules as $L_{\alpha}$ is free. We conclude that $M^{\prime}$ is isomorphic to a direct summand of $M\oplus F$ where $F$ is a finite free $A$-module, and by symmetry $M$ is isomorphic to a direct summand of $M^{\prime}\oplus F^{\prime}$ where $F^{\prime}$ is a finite free $A$-module. If $A$ is complete, it follows from Krull-Schmidt [Leuschke-Wiegand-Cohen-Macaulay-Modules, Corollary 1.10] that $M\oplus P\cong M^{\prime}\oplus P^{\prime}$ for some finite free $A$-modules $P,P^{\prime}$. For a general $A$ the same is true by [Leuschke-Wiegand-Cohen-Macaulay-Modules, Corollary 1.15]. Therefore the Fitting invariants of $M$ and $M^{\prime}$ are the same [stacks, Tag 07ZA].
@@ 1. $(i)$
    The module $C^{a}(L_{A/\mathbf{Z}})$ as in Discussion 11.1 is pseudo-finite-by-flat.
@@ 2. $(ii)$
In fact, Theorem 11.8 holds for all $a<-2$ (cf. Remark 11.4). This is because $R$ can be approximated by countable subrings using [Lyu-elementary-subring] (and [Gor-2-rings, Corollary 3.4]).
@@
Let $\mathbf{Q}$=“$(S_{2})$” and $\mathbf{P}$=“lci.” Let $A\in{{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$. Let $a=-d-2$. Let $\mathfrak{c}(A)$ be the Fitting invariant of $C^{a}(L_{A/\mathbf{Z}})$. Then $\mathfrak{c}(-)$ is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$.
@@
We can find a complete regular local ring $R$ and a surjective ring map $R\to A$. By Theorem 11.8, $\mathfrak{c}(A)$ is well-defined and $0\neq\mathfrak{c}(A)\neq A$. To show $\mathfrak{c}(-)$ is strictly functorial, let $\varphi:A\to B$ be in ${{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$. Apply Discussion 11.2 to
[…]
@@
[…]
where $H=H^{a-1}(L_{B/A})$ and $C=C^{a}(L_{B/A})$. As $\varphi$ is lci and as $a\leq-1$ we have $H=0$ and $C$ flat (in fact projective since $a=-d-2$, cf. proof of Lemma 11.3), so Lemma 11.7$(ii)$ gives $\mathfrak{c}(A)B=\mathfrak{c}(B)$.
@@
We know $M$ is finite flat on the punctured spectrum of $A$ which is lci. If $d\geq 2$, then $0ptA\geq 2$, as $A$ is $(S_{2})$. Therefore the punctured spectrum of $A$ is connected [stacks, Tag 0BLR], so the rank is constant. We may now assume $d=1$.
@@ 12. Local lifting
Let $\mathbf{P}$ be the property “$(S_{k})$$(k\geq 0)$, “Cohen–Macaulay,” “Gorenstein,” or “lci.”
@@
First consider the case $\mathbf{P}$=“$(S_{1})$,” as every Noetherian ring is $(S_{0})$. Let $\mathbf{Q}$ be the trivial property. A strictly functorial $\mathbf{P}$-assignment on $\mathcal{D}_{\mathbf{P}}$ exists, Lemmas 9.2 and 10.1. The assumptions (3) and (4) in Theorem 8.5 are trivial, and we conclude.
Next, consider the case $\mathbf{P}$=“$(S_{2})$.” Let $\mathbf{Q}$=“$(S_{1})$.” A strictly functorial $\mathbf{P}$-assignment on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ exists, Lemmas 9.2 and 10.3. The assumption (4) in Theorem 8.5 is trivial as a domain in $(S_{1})$, whereas (3) follows from the case $\mathbf{P}$=“$(S_{1})$,” and we conclude.
Finally, consider the case $\mathbf{P}$=“$(S_{k})$$(k\geq 3)$, “Cohen–Macaulay,” “Gorenstein,” or “lci.” Let $\mathbf{Q}$=“$(S_{2})$.” A strictly functorial $\mathbf{P}$-assignment on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ exists, Lemma 9.2 and Lemma 10.6, Remark 10.7, Lemma 10.9, and Theorem 11.10. The assumption (3) in Theorem 8.5 follows from the case $\mathbf{P}$=“$(S_{2})$,” and (4) follows from Theorem 7.8 (or [Macaulay-Cesnavi, Corollary 2.14]), and we conclude. ∎
@@
From the case $\mathbf{P}$=Cohen–Macaulay and the same argument as in [Lyu-dual-complex-lift, §8] we get

### v2

@@ 12. A lci-assignment
Let $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$=“lci.” Let $A\in\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$. Intuitively, we want to define $\mathfrak{c}(A)$ to be the Fitting invariant of modules $C_{n}(A/R)$ showing up in [BI23], where $R$ is a regular local ring mapping surjectively to $A$. The flatness of $C_{n}(A/R)$ characterizes lci. However, these modules depend on the choice of $R$ and a projective resolution (in a way that does not change the Fitting invariant, however), and are fragile along ascent (i.e. still involve non-finite modules). We will work with $C_{n}(A/\mathbf{Z})$ instead, which gives the same Fitting invariant. We use standard notations for derived categories, and cohomological conventions for cotangent complexes, as in [Stacks].
Discussion 12.1.
@@
The collection of all such $C^{a}(P^{\bullet})$ is denoted $\mathcal{C}^{a}(L)$. For $X,Y\in\mathcal{C}^{a}(L)$, there exist projective modules $P,Q$ so that $X\oplus P\cong Y\oplus Q$, see [BI23, (7.2)]. We write $C^{a}(L)$ for an unspecified element in $\mathcal{C}^{a}(L)$. The flat and projective dimensions of $C^{a}(L)$ are well-defined.
If $L$ has tor-amplitude in $[a,b]$, then $C^{a}(L)$ is flat. Indeed, let $P^{\bullet}$ represent $L$, then $\tau_{\geq a}(P^{\bullet}\otimes_{A}M)$ represents $L\otimes_{A}^{L}M$ for all $A$-modules $M$, as $L\otimes_{A}^{L}M\in D^{[a,b]}(A)$. Unwinding the definitions, we see $C^{a}(P^{\bullet})\otimes_{R}M=C^{a}(P^{\bullet})\otimes^{L}_{R}M$, as desired. This also tells us if $C^{a}(L)$ has projective dimension $p<\infty$, then $L$ has projective-amplitude in [ap,b][a-p,b].
@@
$\begin{CD}H^{a-1}(L^{\prime\prime})@>{}>{}>C^{a}(P^{\prime\bullet})@>{}>{}>C^{a}(P^{\bullet})@>{}>{}>C^{a}(P^{\prime\prime\bullet})@>{}>{}>0.\end{CD}$
@@
Apply Discussion 12.2 to the triangle
$\begin{CD}L_{A/\mathbf{Z}}@>{}>{}>L_{A/R}@>{}>{}>(L_{R/\mathbf{Z}}\otimes^{L}_{R}A)[1]@>{}>{}>+1,\end{CD}$
@@
$\begin{CD}H@>{}>{}>C^{a}(L_{A/\mathbf{Z}})@>{}>{}>C^{a}(L_{A/R})@>{}>{}>P@>{}>{}>0.\end{CD}$
where $P=C^{a+1}(L_{R/\mathbf{Z}}\otimes^{L}_{R}A)$, $H=H^{a}(L_{R/\mathbf{Z}}\otimes^{L}_{R}A)$. Since $R$ is lci, $L_{R/\mathbf{Z}}$ has tor-amplitude in [1,0][-1,0] [Avr99, (1.2) and (5.1)], so $H=0$. Moreover, every flat $A$-module has projective dimension $\leq\dim A$ [RG71, Seconde partie, Corollaire 3.2.7]. Therefore $L_{R/\mathbf{Z}}$ has projective-amplitude in $\dim A-1,0]$, so $P$ is projective, and we get $C^{a}(L_{A/\mathbf{Z}})\oplus P\cong C^{a}(L_{A/R})$. It remains to observe $C^{a}(L_{A/R})$ is finite up to projective summands, as $L_{A/R}\in D_{Coh}(A)$ [Stacks, Tag 08PZ]. ∎
@@
If $A$ is countable, then we can improve $\dim A-1$ to 2-2, see [RG71, Seconde partie, Corollaire 3.3.2]. We could, if necessary, work extensively with countable rings, via a Löwenheim–Skolem type argument, cf. [Lyu25].
We would like to define the Fitting invariant of $C^{a}(L_{A/\mathbf{Z}})$ to be that of $M$; we will show this is well-defined. Before that, note the following variant of the main theorem of [BI23].
@@
Take the same exact sequence as in the proof of Lemma 12.3. We have $H=0$ and $P$ is flat. Thus $C^{a}(L_{A/\mathbf{Z}})$ is flat if and only $C^{a}(L_{A/R})$ is flat, if and only if $R\to A$ is lci ([BI23, Theorem B] and [Avr99, (1.2)]), if and only if $A$ is lci [Avr99, (5.4) and (5.9)]. ∎
@@ 2. $(ii)$
We have $X=\operatorname{colim}_{\alpha}X_{\alpha}$ as filtered colimits commute with finite limits. Therefore the inclusion $M^{\prime}\subseteq X$ factors through some $X_{\alpha}$ [Stacks, Tag 0G8P]. By [Stacks, Tag 058M] $M^{\prime}\subseteq X$ is pure (i.e. universally injective), so $M^{\prime}\to X_{\alpha}$ is also pure, thus split [Stacks, Tag 058L]. On the other hand $X_{\alpha}\cong M\oplus L_{\alpha}$ as $A$-modules as $L_{\alpha}$ is free. We conclude that $M^{\prime}$ is isomorphic to a direct summand of $M\oplus F$ where $F$ is a finite free $A$-module, and by symmetry $M$ is isomorphic to a direct summand of $M^{\prime}\oplus F^{\prime}$ where $F^{\prime}$ is a finite free $A$-module. If $A$ is complete, it follows from Krull-Schmidt [LW12, Corollary 1.10] that $M\oplus P\cong M^{\prime}\oplus P^{\prime}$ for some finite free $A$-modules $P,P^{\prime}$. For a general $A$ the same is true by [LW12, Corollary 1.15]. Therefore the Fitting invariants of $M$ and $M^{\prime}$ are the same [Stacks, Tag 07ZA].
@@ 1. $(i)$
    The module $C^{a}(L_{A/\mathbf{Z}})$ as in Discussion 12.1 is pseudo-finite-by-flat.
@@ 2. $(ii)$
In fact, Theorem 12.8 holds for all a<2a<-2 (cf. Remark 12.4). This is because $R$ can be approximated by countable subrings using [Lyu25] (and [GM78, Corollary 3.4]).
@@
Let $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$=“lci.” Let $A\in{{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$. Let a=d2a=-d-2. Let $\mathfrak{c}(A)$ be the Fitting invariant of $C^{a}(L_{A/\mathbf{Z}})$. Then $\mathfrak{c}(-)$ is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$.
@@
We can find a complete regular local ring $R$ and a surjective ring map $R\to A$. By Theorem 12.8, $\mathfrak{c}(A)$ is well-defined and $0\neq\mathfrak{c}(A)\neq A$. To show $\mathfrak{c}(-)$ is strictly functorial, let $\varphi:A\to B$ be in $^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$}. Apply Discussion 12.2 to
$\begin{CD}L_{A/\mathbf{Z}}\otimes^{L}_{A}B@>{}>{}>L_{B/\mathbf{Z}}@>{}>{}>L_{B/A}@>{}>{}>+1,\end{CD}$
@@
$\begin{CD}H@>{}>{}>C^{a}(L_{A/\mathbf{Z}})\otimes_{A}B@>{}>{}>C^{a}(L_{B/\mathbf{Z}})@>{}>{}>C@>{}>{}>0\end{CD}$
where $H=H^{a-1}(L_{B/A})$ and $C=C^{a}(L_{B/A})$. As $\varphi$ is lci and as $a\leq-1$ we have $H=0$ and $C$ flat (in fact projective since a=d2a=-d-2, cf. proof of Lemma 12.3), so Lemma 12.7(ii)(ii) gives $\mathfrak{c}(A)B=\mathfrak{c}(B)$.
@@
We know $M$ is finite flat on the punctured spectrum of AA which is lci. If $d\geq 2$, then $0ptA\geq 2$, as $A$ is $(S_{2})$. Therefore the punctured spectrum of $A$ is connected [Stacks, Tag 0BLR], so the rank is constant. We may now assume $d=1$.
@@ 13. Local lifting
Let $\mathbf{P}$ be the property $(S_{k})$$(k\geq 0)$, “Cohen–Macaulay,” “Gorenstein,” or “lci.”
@@
First consider the case $\mathbf{P}$$(S_{1})$,” as every Noetherian ring is $(S_{0})$. Let $\mathbf{Q}$ be the trivial property. A strictly functorial $\mathbf{P}$-assignment on $\mathcal{D}_{\mathbf{P}}$ exists, Lemmas 10.2 and 11.1. The assumptions (3) and (4) in Theorem 9.5 are trivial, and we conclude.
Next, consider the case $\mathbf{P}$$(S_{2})$.” Let $\mathbf{Q}$$(S_{1})$.” A strictly functorial $\mathbf{P}$-assignment on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ exists, Lemmas 10.2 and 11.3. The assumption (4) in Theorem 9.5 is trivial as a domain in $(S_{1})$, whereas (3) follows from the case $\mathbf{P}$$(S_{1})$,” and we conclude.
Finally, consider the case $\mathbf{P}$$(S_{k})$$(k\geq 3)$, “Cohen–Macaulay,” “Gorenstein,” or “lci.” Let $\mathbf{Q}$$(S_{2})$.” A strictly functorial $\mathbf{P}$-assignment on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ exists, Lemma 10.2 and Lemma 11.6, Remark 11.7, Lemma 11.9, and Theorem 12.10. The assumption (3) in Theorem 9.5 follows from the case $\mathbf{P}$$(S_{2})$,” and (4) follows from Theorem 7.8 (or [Čes21, Corollary 2.14]), and we conclude. ∎
@@
From the case $\mathbf{P}$=Cohen–Macaulay and the same argument as in [Lyu, §8] we get